OpenStudy (anonymous):

If ƒ(x ) = x^2 + 1 and g(x ) = 3x + 1, find g(20) + ƒ(6). 74 98 420

OpenStudy (anonymous):

@joyraheb

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

420

OpenStudy (kamibug):

3(20) + 1 = 61 6^2 + 1 = 37 61 + 37 = 98